OpenStudy (anonymous):

2cosx tanx+ √3 tanx =0

OpenStudy (anonymous):

Factor out tan x: tan x(2 cos x + sqrt(3)) = 0 Set each factor equal to 0, you get tan x = 0 and cos x = -sqrt(3)/2 You didnt give an interval for angle x. but tan x = 0 at x = 0, pi, 2pi cos x = - sqrt(3)/2 when x = 5pi/6 and 7pi/6 I used interval [0,2pi]

OpenStudy (anonymous):

nice! yea the interval was between (0, 2pi)

OpenStudy (anonymous):

Thought so....but if it is the OPEN interval (0,2pi), then exclude 0 and 2pi as it is not in the OPEN interval (0,2pi). My answers were based on the closed interval [0,2pi].

OpenStudy (anonymous):

sorry about that yea it was [0,2pi)

OpenStudy (anonymous):

the include 0, exclude 2pi

OpenStudy (anonymous):

Thanks a bunch! :)

OpenStudy (anonymous):

welcome.